A plane flying with the jet stream flew from Los Angeles to Chicago, a distance of 1200 mi, in 4 h. Flying against the jet stream, the plane could fly only 560 mi in the same amount of time.
Find the rate of the plane in calm air and the rate of the wind.

Answer by checkley77(12844)

D=RT
R=D/T
R=1200/4
R=300 MPH. IS THE SPEED WITH THE JET STREAM.
R=560/4
4=140 MPH. IS THE SPEED AGAINST THE WIND.
(300-140)/2=160/2=80 MPH. IS THE SPEED OF THE WIND.
PROOF:
300-80=220 MPH. IS THE SPEED OF THE PLANE IN STILL AIR.
140+80=220 MPH. DITTO.